OpenStudy (anonymous):

Let cos x=1/√10 with x in QIV and find: cot 2x

OpenStudy (anonymous):

cot(2x)=(sin2x)/cos(2x)=2sinxcosx/(2cos^(2)x-1) remember that sinx is -ve in QIV

OpenStudy (anonymous):

\[\cot(2x)=\frac{-2\sqrt{1-\cos^2x} \cos x}{2\cos^2x-1}\]